Explain in detail Lax-wendroff technique. (CFD)
Answer #1

Lax-wendroff technique is mostly used for time and space marching solutions especially to hyperbolic equations. It is a explicit FDM method.

For example we will take a flow field problem governed by a hyperbolic equation.

Time marching solution of a an invicid flow using unsteady Euler equation.  

For demonstrating we will consider an easy Unsteady, 2D inviscid flow

Assuptions :

  1. No body forces
  2. No volumetric heat addition

Non conservative GDE is given as :

Continuity : \frac{\partial\rho }{\partial t}= -\left ( \frac{\partial (\rho u)}{\partial x}+\frac{\partial(\rho V) }{\partial y} \right )== -\left ( \rho \frac{ \partial u}{\partial x}+u\frac{\partial\rho }{\partial x}+\rho \frac{\partial V}{\partial y}+V\frac{\partial \rho }{\partial y} \right )x.momentum: \frac{\partial u}{\partial t}=-\left ( u\frac{\partial u}{\partial x}+v\frac{\partial u}{\partial y}+\frac{1}{\rho }\frac{\partial p}{\partial x} \right )y.momentum: \frac{\partial v}{\partial t}=-\left ( u\frac{\partial V}{\partial x}+V\frac{\partial V}{\partial y}+\frac{1}{\rho }\frac{\partial p}{\partial y} \right )

Energy : \frac{\partial e}{\partial t}=-\left ( u\frac{\partial e}{\partial x}+V\frac{\partial e}{\partial y}+\frac{p}{\rho }\frac{\partial u}{\partial x}+\frac{p}{\rho }\frac{\partial V}{\partial y} \right )

We have to find a numerical soulution using time marching approach

Lax-wendroff technique is based on Taylor series expansion in time.

select any dependent flow variable from \rho, u, V

density at (i,j) at time t+\Delta t

Grid pattern at time

t

t t t tu t 20: + درا -4 زرا- زرا+ دره 2 درا + 2²u 243 دراسل 2(دج) dacat t - زرا+ اد زرا ۔ - راسل + ا+را+نا + + A (As)@y) د- و

Flow VARIABLE ttst CALCULATED list time ذرا۔ i,j it, j Flow VARIABLE Known | -ند t ^x

and at t+\Delta t is given in the picture above picture.

So when you know the fow variable at t, you can find the values at t+\Delta t

Here actually we can directly calculate the flow variable along time direction. No need to solve any set of linear equation. But the method involved is length.

It will give a second order accurate in both space and time

In brief here idea is straight. But procedure is very lengthy.

Hey i thought of giving a typed soulution, but as its a very long one and i was going out of time thats why i switched to hand writing. I know its bit difficult to go through handwriting. Sorry for that. If you have any doubt feel free to comment. I will clarify...If you find this useful dont forget to give me a thumbs up ....Cheers buddy..

Have a nice day